OpenStudy (anonymous):

Determine if triangle XYZ with coordinates X(1, 1), Y(5, 6), and Z(6, 2) is a right triangle. Will medal!

OpenStudy (anonymous):

I need some one to check my answers.

OpenStudy (anonymous):

c=41 a=17 b=26 When I put this into the Pythagorean theorem, it doesn't make sense.

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Mehek14

OpenStudy (dan815):

what are the lengths of the 3 sides?

OpenStudy (anonymous):

For a i got 17, for b 26, and for c 41

OpenStudy (dan815):

those are not hte lengths what u found is a^2, b^2, and c^2

OpenStudy (dan815):

but thats fine

OpenStudy (dan815):

so is a^2+b^2=c^2?

OpenStudy (anonymous):

Should i use the distancxe formula?

OpenStudy (dan815):

yes

OpenStudy (dan815):

draw a picture out to help you

OpenStudy (anonymous):

\[\sqrt{17}+\sqrt{26}=\sqrt{41}\]

OpenStudy (anonymous):

I still need help

OpenStudy (anonymous):

I don't understand what i'm doing